OpenStudy (anonymous):

Where do these two lines intersect? 3x-4y=-25 2x+5y=14

OpenStudy (kira_yamato):

\[\left[\begin{matrix}3 & -4 \\ 2 & 5\end{matrix}\right]\left(\begin{matrix}x \\ y\end{matrix}\right)=\left(\begin{matrix}-25 \\ 14\end{matrix}\right)\] Solve the matrix eqn

OpenStudy (anonymous):